OpenStudy (anonymous):

Four companies donated an average of $135,000 to the Red Cross. If a fifth company donated 170,000 to the Red Cross, what is the average donation for the 5 companies?

OpenStudy (anonymous):

135000*4=540000+170000=71000 710000/5=142000